# Speed, Accuracy, and Complexity

### v1

@@ 3.1.  Characterising Optimal Stopping
Auxillary Payoff Functions
Let us introduce the following auxillary payoff functions: $u_{a}(\alpha,p):=1_{\alpha=a}\cdot(p-\tilde{p})$ and $u_{b}(\alpha,p)=1_{\alpha=b}(\tilde{p}-p)$, where $\tilde{p}$ denotes the belief that renders the decision-maker indifferent between both alternatives, i.e. $\tilde{p}:u(a,\tilde{p})=u(b,\tilde{p})$. The interpretation of these payoff functions is simple: for $\alpha\in A$, $u_{\alpha}$ is such that $\alpha$ is the risky alternative and the other is the safe one, delivering a payoff normalised to zero. This then implies that the relation between beliefs and preference is retained: $u(a,p)\geq u(b,p)\Longleftrightarrow u_{a}(a,p)\geq u_{a}(b,p)\Longleftrightarrow u_{b}(a,p)\geq u_{b}(b,p)\Longleftrightarrow p\geq\tilde{p}$.
We can then define maximised expected payoffs and value functions for both these auxillary payoff functions: For $\alpha\in\{a,b\}$, define $v_{\alpha}(p):=\max_{\hat{\alpha}}u_{\alpha}(\hat{\alpha},p)$ and $V_{\alpha}(p):=\sup_{\tau}\mathbb{E}_{p}[v_{\alpha}(p_{\tau})-(c/\delta)\tau]$.
These will turn out to be quite useful, as the optimal stopping time of the original problem can be characterised in terms of the optimal stopping times of these auxillary problems:
@@
$\operatorname*{arg\,max}_{\tau}\mathbb{E}_{p_{0}}[v(p_{\tau})-c\tau]=\operatorname*{arg\,max}_{\tau}\mathbb{E}_{p_{0}}[v_{a}(p_{\tau})-(c/\delta)\tau]=\operatorname*{arg\,max}_{\tau}\mathbb{E}_{p_{0}}[v_{b}(p_{\tau})-(c/\delta)\tau],$
@@
Relying on these auxillary problems, it is then straightforward to show that the decision-maker’s optimal stopping time is pinned-down by two constant belief thresholds, $\underline{p}$ and $\overline{p}$, such that the decision-maker stops as soon as their belief about $\theta$ reaches either of the two thresholds.
Lemma 2.
@@
This is intuitive, given that $v_{a}$ is increasing in $p$ and $v_{b}$ is decreasing in $p$. Convexity follows from linearity in the measure.
Given this, we can then show that $\underline{p}:=\max\{p\in[0,1]|V_{a}(p)\leq 0\}$ and $\overline{p}:=\min\{p\in[0,1]|V_{b}(p)\leq 0\}$. When $p\leq\underline{p}$, monotonicity of $V_{a}$ implies that $V_{a}(p)=0=u_{a}(b,p)$, and so the decision-maker must stop and choose $b$; and this must be the case in both the auxillary problems. Similarly, when $p\geq\overline{p}$, $V_{b}(p)=0=u_{b}(a,p)$, and so the decision-maker must stop and choose $a$. Hence, the decision-maker has not stopped yet if and only if $p\in(\underline{p},\overline{p})$. Lemma 1 then ties in the argument and yields that the optimal stopping time is characterised by these two thresholds, $\tau^{*}=\inf\{t\geq 0|p_{t}\notin(\underline{p},\overline{p})\}$.
Figure 1 illustrates these lemmas. In panel (a) we have the original maximised expected payoff, $v$, as well as the two auxillary versions $v_{a}$ and $v_{b}$, all of which piece-wise linear functions with a kink at the indifference point $\tilde{p}$. The value functions, $V$, $V_{a}$, and $V_{b}$, are then given on the right panel (b), coinciding with $v$, $v_{a}$, and $v_{b}$, respectively, outside the continuation region $(\underline{p},\overline{p})$.
Refer to caption (a) Refer to caption (b)
Figure 1: Payoff and Value Functions
In showing that the continuation region is characterised by two constant thresholds, Lemma 2 extends this standard result from Shiryaev2007Book to our slightly more general setting in which the payoff difference between the two alternatives is not certain as in FudenbergStrackStrzalecki2018AER. In other words, it is not the fact that $|u(a,\theta)-u(b,\theta)|$ is constant with respect to $\theta$ that underlies the boundary being constant, but rather that $u(a,\theta)-u(b,\theta)$ takes only two possible values.
This observation is valuable as it makes transparent the connection between our model and the drift-diffusion model. In the standard use of the drift-diffusion model, the decision-maker stops as soon as the accumulated evidence reaches a pre-specified threshold (ForstmannRatcliffWagenmakers2016ARPsy), which is equivalent to the decision-maker stopping as soon as their belief about $\theta$ reaches a pre-specified threshold. While these thresholds and the model parameters are typically estimated from choice data, it is understood that the thresholds are endogenous: “[t]he assumption is that payoffs or instructions induce the subject to adjust the positions of the (…) boundaries and so to adjust the amount of information required for a decision” (Ratcliff1978PsyRev). Our focus is exactly on how problem complexity, $\mu/\sigma$, affects the optimal stopping boundaries, and how this in turn affects the speed and accuracy of choices.
@@
Note that $V$ by definition is not a classical solution to the boundary problem since it will necessarily fail to be $\mathcal{C}^{2}$.101010 Note that $V^{\prime\prime}(\overline{p}^{-}),V^{\prime\prime}(\overline{p}^{+})>0=v^{\prime\prime}(\overline{p}^{+}),v^{\prime\prime}(\underline{p}^{-})$. We then obtain the following characterisation of the optimal stopping boundaries. Denoting the log-odds of $p$ by $\ell(p):=\ln(p/(1-p))$, we then have:
@@ 3.2.  Speed and Accuracy
Given the stopping boundaries in log-odds beliefs, $\overline{\ell},\underline{\ell}$, one can translate these to stopping boundaries for the Wiener process, as $\ell_{t}=2(\mu/\sigma^{2})X_{t}$. From there, it is well-known how to obtain closed-form expressions for the expected stopping time and the probability of choosing the best alternative as a function of $\mu/\sigma$. Then, the comparative statics in Theorem 1 arise from a simple, even if cumbersome, application of the implicit function theorem using (1) and (2). Figure 2 exhibits the how speed and accuracy depend on problem complexity and their implied inverse $U$-shaped relationship.
Refer to caption (a) Refer to caption (b) Refer to caption (c)
@@
Theorem 1 implies that easier problems translate into more accurate choices, which, as the proof shows, is due to wider stopping thresholds for beliefs. While wider thresholds entail longer expected stopping times, a higher signal-to-noise ratio entail faster choices. The former effect dominates for more difficult problems, when the signal-to-noise ratio is low, whereas the latter dominates when simpler problems are made simpler.
This result is consistent with existing evidence that, in simple situations, faster choices are more accurate, but in more complicated ones, the opposite is true (e.g. AchtzigerAlos-Ferrer2014MnSc).
We want to highlight two common variants of the model that will not affect this observation. In many instances, the drift is proportional to the payoff difference (e.g. FehrRangel2011JEP). In our model, this would be observationally equivalent to having a drift rate of $\mu\;\delta$ rather than just $\mu$. Theorem 1 would hold just the same in this case. A second variant of the model would feature exponential discounting instead of a flow cost of time; in Appendix B we show that our results are robust: the relation between accuracy and expected stopping time remains non-monotone.
One important implicit assumption in our analysis is that the decision-maker can to some extent recognise the complexity of the problem they face. In this, we differ from previous literature, e.g. FudenbergStrackStrzalecki2018AER, who preclude this possibility. While it may often be the case that one learns about how complicated a particular problem is while trying to solve it, there may be many instances in which there are apparent differences in complexity across problems, and the decision-maker can realise this. In such cases, without further assumptions, one should not in general expect the speed of decision to indicate response accuracy or problem complexity.
@@ 4.  Ability and Effort
Second, we wish to speak to the ambiguity in the literature regarding the relation between ability and response time. In many contexts slower response times are taken to be indicative of greater ability. For instance, it is often found that individuals who commit more time to a problem tend to be more successful or sophisticated (e.g. Rubinstein2016QJE; AgranovCaplinTergiman2015JESA; Esteban-CasanellesGoncalves2020WP; GillProwse2023EJ).
While the studies that relate response time to ability rely on measuring ability from performance independently, one may be tempted to infer ability from response time directly when performance outcomes are ambiguous or not available. Our results suggest that this is not a good idea.
We revisit the model in Section 2 and extend it to allow for the decision-maker to control instantaneous effort $e_{t}\geq 0$, borrowing from MoscariniSmith2001Ecta. Greater effort translates to more informative signals, formally
@@
$\mathbb{E}_{p_{0}}\left[v(p_{\tau})-\int_{0}^{\tau}c(e_{t}/\lambda)\mathop{}\!\mathrm{d}t\right]$
with $p_{t}=\mathbb{E}_{p_{0}}[1_{\theta=a}|X_{t}]$ denoting the decision-maker’s posterior belief, $\mathop{}\!\mathrm{d}X_{t}=\sqrt{e_{t}}\mu s(\theta)\mathop{}\!\mathrm{d}t+\sigma\mathop{}\!\mathrm{d}B_{t}$ and $X_{0}=0$.
@@
We note that Proposition 2 is consistent with recent research in psychology that highlights the importance of taking into account task complexity when interpreting the relation between ability and response time — see e.g. DodonovaDodonov2013Intelligence, GoldhammerNaumannStelterTothRolkeKlieme2014JEducPsy, KyllonenZu2016JIntelligence, and JunghaenelSchneiderOrriensJinLeeKapteynMeijerZelinskiHernandezStone2023JIntelligence. In short, these studies find that, in simple tasks, faster responses are indicative of greater ability, but in complex tasks, the opposite is true. Our simple model can rationalise this nuanced prediction.
@@ 5.  General Costs of Information Acquisition
Specifically, we consider a decision-maker faces uncertainty about an unknown state $\theta\in\Theta$ and can acquire a Blackwell experiment $\pi$ to learn about $\theta$, which, following existing literature (e.g. KamenicaGentzkow2011AER; CaplinDeanLeahy2022JPE; LipnowskiRavid2023WP), we model as a distribution over posterior beliefs, i.e. $\pi\in\Pi:=\Delta(\Delta(\Theta))$ such that $\mathbb{E}_{\pi}[p]=p_{0}$. We denote by $C:\Pi\to\overline{\mathbb{R}}$ the cost of the experiment, which we assume is strictly increasing with respect to the Blackwell order $\geq_{B}$, continuous with respect to the Lévy-Prokhorov metric, strictly convex, and normalised such that a fully uninformative experiment bears zero cost, $C(\delta_{p_{0}})=0$. We model the benefit of information as a function $B:\Pi\to\mathbb{R}$ that is bounded, linear, and strictly increasing with respect to the Blackwell order, which arises as the decision-maker’s expected payoff from choosing an action $a\in A$ in order to maximise their expected payoff given posterior beliefs $p\sim\pi$, i.e. $B(\pi)=\mathbb{E}_{\pi}[\max_{a\in A}\mathbb{E}_{p}[u(a,\theta)]]$.
The decision-maker then chooses an experiment in order to maximise their value of information, $v(\pi;\kappa):=B(\pi)-\kappa C(\pi)$, where $\kappa\geq 0$ is a scaling parameter. We denote the value function, optimal experiment, and cost of the optimal experiment as $V(\kappa):=\sup_{\pi\in\Pi}v(\pi;\kappa)$, $\pi^{*}(\kappa):=\arg\max_{\pi\in\Pi}v(\pi;\kappa)$, and $C^{*}(\kappa):=C(\pi^{*}(\kappa))$.111111 Note that, by Berge maximum theorem, $V$ is continuous, $\pi^{*}$ is single-valued and upper hemicontinuous, and so $C^{*}$ is continuous.
In order to connect this framework to time, we assume that the decision-maker’s expected stopping time is proportional to the total cost of information, i.e. $T^{*}(\kappa):=\mathbb{E}[\tau^{*}]\propto\kappa C^{*}(\kappa)$. The scale $\kappa$ is then interpreted as a measure of problem complexity, scaling the marginal cost of information. In the case of the Wald problem, this is simply reciprocal of the square of the signal-to-noise ratio, $\kappa=(\mu/\sigma)^{-2}$, which extends to a setting with many states (MorrisStrack2019WP).
We then obtain the following result:
@@
The result follows from a simple application of Berge’s maximum theorem.
Single-peakedness in the general case is more difficult to establish, requiring that $-\kappa V^{\prime\prime}(\kappa)/V^{\prime}(\kappa)$ crosses at most once.121212 Note that, since $V$ is convex, by Alexandrov’s theorem, $V^{\prime}$ and $V^{\prime\prime}$ exist a.e., and by an application of the envelope theorem, $\kappa C^{*}(\kappa)=-\kappa V^{\prime}(\kappa)$. Since $\kappa C^{*}(\kappa)$ is non-monotone, $-\kappa V^{\prime\prime}(\kappa)/V^{\prime}(\kappa)$ must cross one at least once; crossing it exactly once is a sufficient and necessary condition for it to be single-peaked. Going beyond this loose characterisation requires additional structure.
@@
Returning to the binary decision problems introduced in Section 2, we suppose $A=\Theta=\{a,b\}$ and $\tilde{p}=1/2$. If one assumes costs of information are posterior-separable (CaplinDeanLeahy2022JPE), i.e. given by $C(\pi)=\mathbb{E}_{\pi}[c(p)]$ for some $c:[0,1]\to\mathbb{R}_{+}$ satisfying $c^{\prime},c^{\prime\prime}>0$, and $c(p_{0})=0$, then $T^{*}(\kappa)$ is single-peaked if and only if $c(p)/c^{\prime}(p)$ is quasiconcave on $[1/2,1]$.

### v2

@@ 3.1.  Characterising Optimal Stopping
Auxiliary Payoff Functions
Let us introduce the following auxiliary payoff functions: $u_{a}(\alpha,p):=1_{\alpha=a}\cdot(p-\tilde{p})$ and $u_{b}(\alpha,p)=1_{\alpha=b}(\tilde{p}-p)$, where $\tilde{p}$ denotes the belief that renders the decision-maker indifferent between both alternatives, i.e. $\tilde{p}:u(a,\tilde{p})=u(b,\tilde{p})$. The interpretation of these payoff functions is simple: for $\alpha\in A$, $u_{\alpha}$ is such that $\alpha$ is the risky alternative and the other is the safe one, delivering a payoff normalised to zero. This then implies that the relation between beliefs and preference is retained: $u(a,p)\geq u(b,p)\Longleftrightarrow u_{a}(a,p)\geq u_{a}(b,p)\Longleftrightarrow u_{b}(a,p)\geq u_{b}(b,p)\Longleftrightarrow p\geq\tilde{p}$.
We can then define maximised expected payoffs and value functions for both these auxiliary payoff functions: For $\alpha\in\{a,b\}$, define $v_{\alpha}(p):=\max_{\hat{\alpha}}u_{\alpha}(\hat{\alpha},p)$ and $V_{\alpha}(p):=\sup_{\tau}\mathbb{E}[v_{\alpha}(p_{\tau})-(c/\delta)\tau]$, where $p=\mathbb{P}(\theta=a)$.
These will turn out to be quite useful, as the optimal stopping time of the original problem can be characterised in terms of the optimal stopping times of these auxiliary problems:
@@
$\operatorname*{arg\,max}_{\tau}\mathbb{E}[v(p_{\tau})-c\tau]=\operatorname*{arg\,max}_{\tau}\mathbb{E}[v_{a}(p_{\tau})-(c/\delta)\tau]=\operatorname*{arg\,max}_{\tau}\mathbb{E}[v_{b}(p_{\tau})-(c/\delta)\tau],$
@@
Relying on these auxiliary problems, it is then straightforward to show that the decision-maker’s optimal stopping time is pinned-down by two constant belief thresholds, $\underline{p}$ and $\overline{p}$, such that the decision-maker stops as soon as their belief about $\theta$ reaches either of the two thresholds.
@@
This is intuitive, given that $v_{a}$ is increasing in $p$ and $v_{b}$ is decreasing in $p$. Convexity, a property that will be useful later on, follows from linearity in the measure.
Given this, we can then show that
Lemma 3.
In particular, we have that $\underline{p}:=\max\{p\in[0,1]|V_{a}(p)\leq 0\}$ and $\overline{p}:=\min\{p\in[0,1]|V_{b}(p)\leq 0\}$. When $p\leq\underline{p}$, monotonicity of $V_{a}$ implies that $V_{a}(p)=0=u_{a}(b,p)$, and so the decision-maker must stop and choose $b$; and this must be the case in both the auxiliary problems. Similarly, when $p\geq\overline{p}$, $V_{b}(p)=0=u_{b}(a,p)$, and so the decision-maker must stop and choose $a$. Hence, the decision-maker has not stopped yet if and only if $p\in(\underline{p},\overline{p})$. Lemma 1 then ties in the argument and yields that the optimal stopping time is characterised by these two thresholds, $\tau^{*}=\inf\{t\geq 0|p_{t}\notin(\underline{p},\overline{p})\}$.
Figure 1 illustrates these lemmas. In panel (a) we have the original maximised expected payoff, $v$, as well as the two auxiliary versions $v_{a}$ and $v_{b}$, all of which piece-wise linear functions with a kink at the indifference point $\tilde{p}$. The value functions, $V$, $V_{a}$, and $V_{b}$, are then given on the right panel (b), coinciding with $v$, $v_{a}$, and $v_{b}$, respectively, outside the continuation region $(\underline{p},\overline{p})$.
Refer to caption (a) Refer to caption (b) Figure 1: Payoff and Value Functions
In showing that the continuation region is characterised by two constant thresholds, Lemma 3 extends this standard result from Shiryaev (2007) to our slightly more general setting in which the payoff difference between the two alternatives is not certain as in Fudenberg et al. (2018). In other words, it is not the fact that $|u(a,\theta)-u(b,\theta)|$ is constant with respect to $\theta$ that underlies the boundary being constant, but rather that $u(a,\theta)-u(b,\theta)$ takes only two possible values.
This observation is valuable as it makes transparent the connection between our model and the drift-diffusion model. In the standard use of the drift-diffusion model, the decision-maker stops as soon as the accumulated evidence reaches a pre-specified threshold (Forstmann et al., 2016), which is equivalent to the decision-maker stopping as soon as their belief about $\theta$ reaches a pre-specified threshold. While these thresholds and the model parameters are typically estimated from choice data, it is understood that the thresholds are endogenous: “[t]he assumption is that payoffs or instructions induce the subject to adjust the positions of the (…) boundaries and so to adjust the amount of information required for a decision” (Ratcliff, 1978). Our focus is exactly on how problem complexity, $\mu/\sigma$, affects the optimal stopping boundaries, and how this in turn affects the speed and accuracy of choices.
@@
Note that $V$ by definition is not a classical solution to the boundary problem since it will necessarily fail to be $\mathcal{C}^{2}$.151515 Note that $V^{\prime\prime}(\overline{p}^{-}),V^{\prime\prime}(\overline{p}^{+})>0=v^{\prime\prime}(\overline{p}^{+}),v^{\prime\prime}(\underline{p}^{-})$.
We then obtain a sharp characterisation of the optimal stopping boundaries. To state it, let the log-odds of $p$ be denoted by $\ell(p):=\ln(p/(1-p))$. The log-odds representation will be useful, since the log-odds of the posterior beliefs at time $t$ is linear in $X_{t}$:
$\ell_{t}=\ln\left(\frac{\mathbb{P}(\theta=a|X_{t})}{\mathbb{P}(\theta=b|X_{t})}\right)=\ln\left(\frac{\phi((X_{t}-\mu\,t)/\sqrt{t}\sigma)}{\phi((X_{t}+\mu\,t)/\sqrt{t}\sigma)}\right)=2(\mu/\sigma^{2})X_{t}.$
We then have:
@@ 3.2.  Speed and Accuracy
Given the stopping boundaries in log-odds beliefs, $\overline{\ell},\underline{\ell}$, one can translate these to stopping boundaries for the Wiener process, as $\ell_{t}=2(\mu/\sigma^{2})X_{t}$. From there, it is well-known how to obtain closed-form expressions for the expected stopping time and the probability of choosing the best alternative as a function of $\mu/\sigma$. Then, the comparative statics in Theorem 1 arise from a simple, even if cumbersome, application of the implicit function theorem using (1) and (2). Figure 2 exhibits how speed and accuracy depend on problem complexity and their implied inverse $U$-shaped relationship.
@@
While, at first glance, that accuracy is monotone in problem complexity may not seem particularly surprising, it is important to note that the result in Theorem 1 is independent of ex-post relative preference intensity. In particular, it might be more important for the decision-maker to choose $\alpha=\theta$ when $\theta=a$ than when $\theta=b$. For instance, when alternative $a$ is better it is much better than $b$, but when it is worse it is only mildly worse — something that would be captured by a small value of the indifference parameter $\tilde{p}$. Our results show that such considerations regarding relative preference intensity do not matter: the probability with which the decision-maker chooses the ex-post optimal alternative is monotone in problem complexity.
Theorem 1 means that easier problems translate into more accurate choices, which, as the proof shows, is due to wider stopping thresholds for beliefs. While wider thresholds entail longer expected stopping times, a higher signal-to-noise ratio entail faster choices. The former effect dominates for more difficult problems, when the signal-to-noise ratio is low, whereas the latter dominates when simpler problems are made simpler.
Theorem 1 also implies that easier problems translate into more accurate choices, which, as the proof shows, is due to wider stopping thresholds for beliefs. While wider thresholds entail longer expected stopping times, a higher signal-to-noise ratio entail faster choices. The former effect dominates for more difficult problems, when the signal-to-noise ratio is low, whereas the latter dominates when simpler problems are made simpler. This result is consistent with existing evidence that, in simple situations, faster choices are more accurate, but in more complicated ones, the opposite is true (e.g. Onken et al., 1985; Achtziger and Alós-Ferrer, 2013).
We want to highlight two common variants of the model that will not affect this observation. In many instances, the drift is proportional to the payoff difference (e.g. Fehr and Rangel, 2011). In our model, this would be observationally equivalent to having a drift rate of $\mu\;\delta$ rather than just $\mu$. Theorem 1 would hold just the same in this case. A second variant of the model would feature exponential discounting instead of a flow cost of time; in Appendix B we show that our results are robust: the relation between accuracy and expected stopping time remains non-monotone.
One important implicit assumption in our analysis is that the decision-maker can to some extent recognise the complexity of the problem they face. In this, we differ from previous literature, e.g. Fudenberg et al. (2018), who preclude this possibility. While it may often be the case that one learns about how complicated a particular problem is while trying to solve it, there may be many instances in which there are apparent differences in complexity across problems, and the decision-maker can realise this. In such cases, without further assumptions, one should not in general expect the speed of decision to indicate response accuracy or problem complexity.
@@ 4.  Ability and Effort
Second, we wish to speak to the ambiguity in the literature regarding the relation between ability and response time. In many contexts slower response times are taken to be indicative of greater ability or sophistication. For instance, more financially literate individuals will by definition be at a better position for understanding information about alternative financial assets (e.g. Darriet et al., 2020). Also in strategic settings, it is often found that individuals who commit more time to a problem tend to be more successful or sophisticated (e.g. Rubinstein, 2016; Agranov et al., 2015; Alós-Ferrer and Bruckenmaier, 2021; Esteban-Casanelles and Gonçalves, 2022; Gill and Prowse, 2023).
While the studies that relate response time to ability rely on measuring ability from performance independently, one may be tempted to infer ability from response time directly when performance outcomes are ambiguous or not available. Anecdotes and introspection again caution against this. For example, students who return their exam papers early would believe that spending more time and effort would not benefit them, but this could indicate either strong performance or a lack of progress. Our results corroborate this intuitive ambiguity.
We revisit the model in Section 2 and extend it to allow for the decision-maker to control instantaneous effort $e_{t}\geq 0$, borrowing from Moscarini and Smith (2001). Greater effort translates to more informative signals, formally
@@
$\mathbb{E}\left[v(p_{\tau})-\int_{0}^{\tau}c(e_{t}/\lambda)\mathop{}\!\mathrm{d}t\right]$
with $p_{t}=\mathbb{E}[1_{\theta=a}|X_{t}]$ denoting the decision-maker’s posterior belief, $\mathop{}\!\mathrm{d}X_{t}=\sqrt{e_{t}}\mu\,s(\theta)\mathop{}\!\mathrm{d}t+\sigma\mathop{}\!\mathrm{d}B_{t}$ and $X_{0}=0$.
@@
We note that Proposition 2 is consistent with recent research in psychology that highlights the importance of taking into account task complexity when interpreting the relation between ability and response time — see e.g. Dodonova and Dodonov (2013), Goldhammer et al. (2014), Kyllonen and Zu (2016), and Junghaenel et al. (2023). In short, these studies find that, in simple tasks, faster responses are indicative of greater ability, but in complex tasks, the opposite is true. Our simple model can rationalise this nuanced prediction.
@@ 5.  General Costs of Information Acquisition
Specifically, we consider a decision-maker faces uncertainty about an unknown parameter or event $\theta\in\Theta$ and can acquire a Blackwell experiment $\pi$ to learn about $\theta$, which, following existing literature (e.g. Kamenica and Gentzkow, 2011; Caplin et al., 2022; Lipnowski and Ravid, 2023), we model as a distribution over posterior beliefs, i.e. $\pi\in\Pi:=\Delta(\Delta(\Theta))$ such that $\mathbb{E}_{\pi}[p]=p_{0}$. We denote by $C:\Pi\to\overline{\mathbb{R}}$ the cost of the experiment, which we assume is strictly increasing with respect to the Blackwell order $\geq_{B}$, continuous with respect to the Lévy-Prokhorov metric, strictly convex, and normalised such that a fully uninformative experiment bears zero cost, $C(\delta_{p_{0}})=0$. We model the benefit of information as a function $B:\Pi\to\mathbb{R}$ that is bounded, linear, and strictly increasing with respect to the Blackwell order, which arises as the decision-maker’s expected payoff from choosing an action $a\in A$ in order to maximise their expected payoff given posterior beliefs $p\sim\pi$, i.e. $B(\pi)=\mathbb{E}_{\pi}[\max_{a\in A}\mathbb{E}_{p}[u(a,\theta)]]$.
The decision-maker then chooses an experiment in order to maximise their value of information, $v(\pi;\kappa):=B(\pi)-\kappa C(\pi)$, where $\kappa\geq 0$ is a scaling parameter. We denote the value function, optimal experiment, and cost of the optimal experiment as $V(\kappa):=\sup_{\pi\in\Pi}v(\pi;\kappa)$, $\pi^{*}(\kappa):=\arg\max_{\pi\in\Pi}v(\pi;\kappa)$, and $C^{*}(\kappa):=C(\pi^{*}(\kappa))$.161616 Note that, by Berge maximum theorem, $V$ is continuous, $\pi^{*}$ is single-valued and upper hemicontinuous, and so $C^{*}$ is continuous.
In order to connect this framework to time, we assume that the decision-maker’s expected stopping time is proportional to the total cost of information, i.e. $T^{*}(\kappa):=\mathbb{E}[\tau^{*}]\propto\kappa C^{*}(\kappa)$. The scale $\kappa$ is then interpreted as a measure of problem complexity, scaling the marginal cost of information. In the case of the Wald problem, this is simply reciprocal of the square of the signal-to-noise ratio, $\kappa=(\mu/\sigma)^{-2}$, which extends to a setting with many states (Morris and Strack, 2019).171717 See Pomatto et al. (2023) for an axiomatic characterisation of such cost function.
We then obtain the following result from a simple application of Berge’s maximum theorem:
@@
Single-peakedness in the general case is more difficult to establish, requiring that $-\kappa V^{\prime\prime}(\kappa)/V^{\prime}(\kappa)$ crosses at most once.181818 Note that, since $V$ is convex, by Alexandrov’s theorem, $V^{\prime}$ and $V^{\prime\prime}$ exist a.e., and by an application of the envelope theorem, $\kappa C^{*}(\kappa)=-\kappa V^{\prime}(\kappa)$. Since $\kappa C^{*}(\kappa)$ is non-monotone, $-\kappa V^{\prime\prime}(\kappa)/V^{\prime}(\kappa)$ must cross one at least once; crossing it exactly once is a sufficient and necessary condition for it to be single-peaked. Going beyond this loose characterisation requires additional structure.
@@
Returning to the binary decision problems introduced in Section 2, we suppose $A=\Theta=\{a,b\}$ and $\tilde{p}=1/2$. If one assumes costs of information are posterior-separable (Caplin et al., 2022), i.e. given by $C(\pi)=\mathbb{E}_{\pi}[c(p)]$ for some $c:[0,1]\to\mathbb{R}_{+}$ satisfying $c^{\prime},c^{\prime\prime}>0$, and $c(p_{0})=0$, then $T^{*}(\kappa)$ is single-peaked if and only if $c(p)/c^{\prime}(p)$ is quasiconcave on $[1/2,1]$.